# Survey in Characterization of Semantic Change

### v2

@@ 5 Formalization
In statistics (and machine learning), the problem of drift distribution is formally defined Gama et al., (2014). However, in linguistics, the definition of word’s meaning (from a “semasiological” perspective) still relies on ideas such as ‘concepts’ and ‘sentiments’ without providing objective definitions for this terms Koch, (2016); Traugott, (2017); Blank et al., (2003), making a formal definition of drift distribution more challenging. This vague notion of change complicates the implementation of procedural methodologies to analyse, compare, and evaluate LSC. In this section, we based our formalisation on the definitions coming from the statistical domain.
First, we assume that for a set of corpora of text $T$ there is a universe of senses $S_{T}$, that holds all the senses for that corpus. Given an alphabet $\alpha$, we define $\mathcal{S}$ as a function that takes a string of the closure $\alpha$ where $\alpha^{*}=V$ (the vocabulary) and returns a subset of $S_{T}$. For instance, given a word $w\in V$ and a corpus $t\in T$ , the subset of senses $S(w,t)$ for the universe of senses ($S_{t}$) of the word $w$ in this corpus is the following:
@@
$\begin{split}\mathcal{S}(a,x)&\neq\mathcal{S}(b,y)\iff\\ (\mathcal{S}(a,x)-\mathcal{S}(b,y)\neq\emptyset)&\lor(\mathcal{S}(b,y)-\mathcal{S}(a,x)\neq\emptyset)\end{split}$ (2)
@@ 5.1 Semantic Change
We define that $w$ had a semantic change ($\mathcal{C}$), if the set of senses, given by the function $\mathcal{S}$, evaluated in the corpus $t_{1}$ is different from the set of senses in $t_{2}$, given that $t_{1}\neq t_{2}$,
$\mathcal{C}(x,t_{1},t_{2})=\begin{cases}\text{True}&\mathcal{S}(x,t_{1})\neq\mathcal{S}(x,t_{2})\\ \text{False}&\text{otherwise.}\end{cases}$ (3)
@@ 5.2 Change in Dimension
This definition is compatible with Koch, (2016) concept of innovative/reductive meaning change, in which words acquire new senses (sememes) by being used with different meaning. However, it differs for the more general broadening notion, where words that undergo broadening may gain extension, meaning that they are used in a broader sense than initially, but they may also lose intention, which is the precision of the word’s meaning, making it more vague.
@@ 5.3 Change in Relation
We define that a word $w$ had changed in relation, if the aggregated sum of the relational similarity ($\mathit{l}:S\times S\rightarrow\mathbf{R}$) between senses in $\mathcal{R}(w,t_{1})$ changed for $\mathcal{R}(w,t_{2})$. Where $\mathit{l}$ takes two senses $s_{i},s_{j}$ and returns a real value, that correspond to the material or abstract relation between these senses (greater the value, more related the senses are), i.e.,
@@
Given a word $w$, we define that it increased in relation with respect to corpus $t_{1},t_{2}$, if $\mathcal{R}(w,t_{1})<\mathcal{R}(w,t_{2})$, and decreased if the opposite happens. This definition relies on the notion of material or abstract similarity between senses that we still need to clarify for figures of language, for instance, metaphor and metonymy.
The phenomenon of metaphorization or metonymization occurs when a word takes on a new meaning that, to some extent, inherits qualities from its original meaning through a figurative relationship the speaker aims to convey Lakoff and Johnson, (2008). The original meaning is invoked by establishing a tangible or abstract connection with the initial sense, where the listener is able to solve the non-literal source concept.
A material relation is built when we recur to some cultural association established between the concepts (source and target) that are interleaved under a shared context (a frame) Koch, (2016), e.g., in the sentence “The proposition made by Washington was rejected by the other countries”, we are replacing the location of the institution by the institution itself. Other forms of metonymy use cause and effect, contiguity, instrumentality, possession, etc [^3] Littlemore and Tagg, (2018) group the forms of metonymy as ‘referential,’ ‘propositional,’ ‘illocutionary,’ and ‘situational.’. So, for a reader be able to solve this semantics, is mandatory a cultural knowledge of the type of link between the source and target concepts that the speaker is trying to evoke.
On the other hand, an abstract relation is achieved by associating properties that evoke sensory experiences, such as emotions, shapes, dimensions, positions, and more Lakoff and Johnson, (2008). In this case, the literal sense is elucidated by delving into the figurative properties that the listener can grasp within the sense itself.
The theory on how metonymies and metaphors are defined is ambiguous, where generally relying on subjective perception, such as personal experience, or definitions that overlap with other figurative modes. For example, Choi et al., (2021) defines metaphor as “literal meaning of a word is different from its contextual meaning”, which is also the case for metonymy. Many authors in linguistic try to disambiguate the relation relying on imprecise definitions Burke, (1969), and as discussed by Koch, (2016) there’s no precise rule to define a metonymy. In Schlechtweg et al., (2017) the authors only measure the entropy of a word as a means to capture metaphorization, it does not provide evidences that this measure does not vary for the case of metonymization or a new sense not related to the previous senses  [^4] For example, the word ‘apple’ used in a technology context increases entropy but is not a metaphorization process, but a new sense..
@@ 5.4 Change in Orientation
We define that a word $w$ had changed in orientation, if for some notion of orientation $\mathfrak{f}$ (feeling), there is a mapping from a word $w$ in a corpora $t$ to an orientation value, for example, in case of positive or negative feeling,
$\mathfrak{f}:V\times T\rightarrow\{-1,+1\},$ (7)
@@ 5.5 Semantic Change Identification
Consider a function $f_{\mathcal{C}}$ that takes two representations $\mathfrak{R}$ (e.g., embeddings, graphs, etc.) of the word $w$ in $t_{1},t_{2}$. We define the problem of semantic change identification as assigning a label (changed or not changed) to this word,
$f_{\mathcal{C}}(\mathfrak{R}(w,t_{1}),\mathfrak{R}(w,t_{2}))\rightarrow\{0,1\},$ (9)
which tries to approximate the set of words $\bar{W}_{t_{1},t_{2}}$ where we can measure some notion of error, e.g., accuracy.
@@ 5.6 Semantic Change Characterization
We define semantic change characterization as a classification problem where we need to assign identified semantic changes to categories, such as $f_{\mathcal{D}},f_{\mathcal{R}},f_{\mathcal{O}}$,
@@ 6.1 Dimension
The majority of the studies found during this survey propose methods and tools to characterize how words change in dimension, while the other poles were less studied. Change in dimension seems to be easier to detect than the other categories because it does not rely on subjective interpretation, such as relation or feeling, but mainly on countable quantity (see Section 2.2). Works along this axis often rely on word sense induction, a research area with well developed tooling and many data sources, such as dictionaries and synsets, used to implement supervised learning.
A major characteristic that differentiate the approaches is the word representation method. We identified four main types of word representation methods, used to study dimension changes in LSC. These methods have different trade-offs in implementation cost, representation power, and interpretability. It is difficult to compare the effectiveness of each method for this state-of-the-art, given the qualitative and exploratory nature of the reported results, the use of different corpora (not always available), and the need for external knowledge to acquire meaning (difficult to reproduce). However, the analysis reveals that word frequency and embeddings-based methods produce better results to identify neologism (new words), while topics and graph-based approaches are more effective at detecting sememes (new senses) over time. A limitation for the approaches implemented based on dictionaries is that the meaning of new words are not always in the dictionaries because there is a temporal delay between the first occurrences of these neologisms and their publication in dictionaries. Another limitation comes with the presence, in dictionaries, of metaphorical senses at the same level as the literal ones, increasing the complexity to make the distinction between ‘dimension’ and ‘relation’. Some authors prefer to group these two poles into one, as in Tang et al., (2013) . The limitation for topics- and graph-based approaches is that the interpretation of the meaning of a topic (or a branch of a graph) is not always clear and can require the intervention of an expert to, for instance, to regroup topics/branches with the same meaning and/or to map these groups to word’s senses.
@@ 6.2 Relation
Outside of this LSC survey, there is a vast body of work proposing methods and tools to predict if a word in a sentence has a metaphorical sense Heintz et al., (2013); Shutova and Sun, (2013); Hovy et al., (2013). These approaches analyze texts, but they do not consider the temporal information, thus they cannot express if a word suffers from metaphorization over time. Some works claim to detect metaphorization of words (e.g., Tredici et al., (2016) ), but they lay on the general hypothesis that changing sense of words implies in a metaphorical change, which still needs to be demonstrated.
For metonymization, we could not find studies that performed metonymy identification in a comparative setting. However, related recent studies investigate the usage of metonymy within different methods to build the material relation identification Nastase et al., (2012); Soares et al., (2019); Mathews and Strube, (2021). We expect that future investigations can extend these approaches to allow detecting metonymization of words.
@@
Tang et al., (2013) mainly rely on dictionaries of heuristics to infer metaphoricity, which served for qualitative analysis, but still require more extensive validation. A single work used ground-truth annotations for metaphors Schlechtweg et al., (2017), but did not reported a final precision of the proposed methodology.
@@ 6.3 Orientation
Although there is a huge body of work proposing solutions for analysing sentiment at sentence level, fewer of them try to infer sentiments at a word-level Guo et al., (2021), and in the word-level, only a small fraction is performed on a comparative manner, as we accessed in this survey.
An important outcome of this survey is that there is an harmonization on how change in feeling are defined and applied. However, we could not observe that there is also an harmonization on how the error analysis for the feeling function is implemented. Some works adopt binary orientation function (positive, negative), while other multidimensional orientations (valence, arousal, dominance) Buechel et al., (2016). There is still further study to analyze which function ($\mathfrak{f}$) is more suitable for characterizing change.
From the many approaches analysed in this survey, one can observed some converging points. This is the case for instance for the difficulty to evaluate and compare the obtained results because, for the majority of the tasks, no gold standard exists. This problem has been raised in many papers like Schlechtweg et al., (2017), but some initiatives to create reference datasets are under developing.

### v3

@@ 5 Formalization
In computer science, formalizing a problem is an essential step. It allows us to reason about the problem using computational concepts and directly design algorithms to solve it. This process is particularly crucial in research fields. A well-defined formalization ensures that researchers are working with the same underlying assumptions and terminology. This shared ground facilitates the development and comparison of different methodologies. In our review, we observed that while the analyzed papers conceptually addressed the same problem, the lack of a common formalization made it difficult to compare them effectively. To address this gap, this section proposes a formalization based on the reviewed studies and insights gleaned from linguistic literature to guide future research in this area.
However, it’s important to acknowledge the limitations of purely computational approaches. Linguistic literature, exemplified by Campbell, (2013, p. 221-246), explores semantic change through historical and socio-cultural lenses, examining how societal events influence word evolution. Replicating this approach entirely through computational analysis remains an open challenge. Therefore, we focus on formalizing methods based solely on semasiological comparative analysis, which compares the meaning of words across time without relying on additional knowledge of the cultural context in which those meanings were produced.
In statistics (and machine learning), the problem of drift distribution is formally defined (Gama et al.,, 2014). However, in linguistics, the definition of word’s meaning (from a “semasiological” perspective) still relies on ideas such as ‘concepts’ and ‘sentiments’ without providing objective definitions for this terms (Koch,, 2016; Traugott,, 2017; Blank et al.,, 2003), making a formal definition of drift distribution more challenging. This vague notion of change complicates the implementation of procedural methodologies to analyse, compare, and evaluate LSC.
We build our formalization by posing the meaning of a word as as a set of glosses, inspired by Hock and Joseph, (2019, p. 194), and define all constructions in set theory. First, we assume that for a set of corpora of text $T$ there is a universe of senses $S_{T}$, that holds all the senses for that corpus. Given an alphabet $\alpha$, we define $\mathcal{S}$ as a function that takes a string of the closure $\alpha$ where $\alpha^{*}=V$ (the vocabulary) and returns a subset of $S_{T}$. For instance, given a word $w\in V$ and a corpus $t\in T$ , the subset of senses $S(w,t)$ for the universe of senses ($S_{t}$) of the word $w$ in this corpus is the following:
@@
$\begin{split}\mathcal{S}(a,x)&\neq\mathcal{S}(b,y)\iff\\ (\mathcal{S}(a,x)-\mathcal{S}(b,y)\neq\emptyset)&\lor(\mathcal{S}(b,y)-\mathcal{S}(a,x)\neq\emptyset),\end{split}$ (2)
or in simple terms, they are different if the sets of senses from corpus $t_{1}$ and $t_{2}$ have different elements.
@@ 5.1 Semantic Change
Koch, (2016, p. 23) presents a definition of meaning change that combines semasiological and onomasiological perspective. In Figure 4, ‘belly’ was initially associated with the source concept ‘bag’ (M1), later it acquired the meaning of the target concept ‘body part between the breast and the thighs’ (M2). However, in the field of computational linguistics, we primarily focus on the semasiological view. Here, our interest lies solely in the word itself and the evolution of its meaning, rather than delving into the origins of the concepts it represents.
Figure 4: Definition of semantic change adapted from Koch, (2016, p. 23,25).
The process of meaning change is depicted on the right side of the figure. At time T1, the word ”belly” possessed only the meaning M1. Subsequently, at T2, it began to be used with the new meaning M2 as well. This usage of M2 increased over time (T3-T5) until it eventually became the dominant meaning (T6). Finally, at T7, the original meaning M1 was completely lost.
To formalize this definition, we say that a word $w$ had a semantic change ($\mathcal{C}$), if the set of senses, given by the function $\mathcal{S}$, evaluated in the corpus $t_{1}$ is different from the set of senses in $t_{2}$, given that $t_{1}\neq t_{2}$,
$\mathcal{C}(w,t_{1},t_{2})=\begin{cases}\text{True}&\mathcal{S}(w,t_{1})\neq\mathcal{S}(w,t_{2})\\ \text{False}&\text{otherwise.}\end{cases}$ (3)
@@ 5.2 Change in Dimension
This definition is compatible with Koch, (2016, p. 24-26) concept of innovative/reductive meaning change, in which words acquire new senses (sememes) by being used with different meaning. However, it differs for the more general broadening notion, where words that undergo broadening may gain extension, meaning that they are used in a broader sense than initially, but they may also lose intention, which is the precision of the word’s meaning, making it more vague.
@@ 5.3 Change in Relation
The phenomenon of metaphorization or metonymization occurs when a word takes on a new meaning that, to some extent, inherits qualities from its original meaning through a figurative relationship the speaker aims to convey (Lakoff and Johnson,, 2008, p. 3-6). The original meaning is invoked by establishing a material or abstract relation with the initial sense, where the listener is able to solve the non-literal target concept. In the realm of human cognition, material relations and abstract relations are distinguished by the type of association formed between concepts. Material relations are established through the creation of contiguity associations in the physical world. This means that the relationship between the target and source concepts is built upon their physical proximity or interaction. In contrast, abstract relations are formed through similarity associations within the conceptual world. Here, the connection between concepts arises from shared characteristics or underlying principles, independent of any physical manifestation.
A material relation is built when we recur to some cultural association established between the concepts (source and target) that are interleaved under a shared context (a frame) (Koch,, 2016, p. 42-51), e.g., in the sentence “The proposition made by Washington was rejected by the other countries”, we are replacing the location of the institution by the institution itself. Other forms of metonymy use cause and effect, contiguity, instrumentality, possession, etc [^4] Littlemore and Tagg, (2018) group the forms of metonymy as ‘referential,’ ‘propositional,’ ‘illocutionary,’ and ‘situational.’.
On the other hand, an abstract relation is achieved by associating properties that evoke sensory experiences, such as emotions, shapes, dimensions, positions, and more (Lakoff and Johnson,, 2008, p. 56-60). For example, “The arm of the chair is broken,” in this sentence the listener imagine the chair as a human being and associate body parts to relative positions of the chair. In this case, the literal sense is elucidated by examining the figurative properties that the listener can grasp within the sense itself.
The theory on how metonymies and metaphors are defined is informal, where generally relying on subjective perception, such as personal experience, or definitions that overlap with other figurative modes. For example, Choi et al., (2021) defines metaphor as “literal meaning of a word is different from its contextual meaning”, which is also the case for metonymy. Many authors in linguistic try to disambiguate the relation relying on imprecise definitions (Campbell,, 2013, p.224-225), and as discussed by Koch, (2016, p. 42-51) there’s no precise rule to define a metonymy. In Schlechtweg et al., (2017) the authors only measure the entropy of a word as a means to capture metaphorization, it does not provide evidences that this measure does not vary for the case of metonymization or a new sense not related to the previous senses  [^5] For example, the word ‘apple’ used in a technology context increases entropy but is not a metaphorization process, but a new sense..
Given the above considerations, we define that a word $w$ had changed in relation, if the aggregated sum of the relational similarity ($\mathit{l}:S\times S\rightarrow\mathbf{R}$) between senses in $\mathcal{R}(w,t_{1})$ changed for $\mathcal{R}(w,t_{2})$. Where $\mathit{l}$ takes two senses $s_{i},s_{j}$ and returns a real value, that correspond to the material or abstract relation between these senses (greater the value, more related the senses are), i.e.,
@@
Given a word $w$, we define that it increased in relation with respect to corpus $t_{1},t_{2}$, if $\mathcal{R}(w,t_{1})<\mathcal{R}(w,t_{2})$, and decreased if the opposite happens.
@@ 5.4 Change in Orientation
The change in word’s orientation (positive or negative connotation) reflects the amelioration or pejoration in the usage of a word from one corpus to another. This is measured by analyzing how the collective perception of a word’s meaning evolves within a specific context (Campbell,, 2013, p. 227-229). A classical example is the word ‘terrific’. Originally, ‘terrific’ meant ‘terrible’, carrying a negative connotation. Then, its meaning shifted to ‘formidable’, signifying something positive and impressive.
Buechel et al., (2016) uses a Valence-Arousal-Dominance setting, where the valence axis represent positive and negative sentiment, but it also extends the emotion perception to more dimensions, arousal and dominance. While this in-depth analysis has its value, it is not explored in the linguistic literature and the majority of papers in sentiment analysis (Cui et al.,, 2023), so we restrict our formalism to the more general polarity setting(positive/negative), to be aligned with the current research and applications (Susanto et al.,, 2020).
We define that a word $w$ had changed in orientation, if for some notion of orientation $\mathfrak{f}$ (feeling), there is a mapping from a word $w$ in a corpora $t$ to an orientation value, for example, in case of positive, neutral, or negative feeling,
$\mathfrak{f}:V\times T\rightarrow\{-1,0,+1\},$ (7)
@@ 5.5 Semantic Change Identification
Consider a function $f_{\mathcal{C}}$ that takes two representations $\mathfrak{R}$ (e.g., embeddings, graphs, etc.) of the word $w$ in corpora $t_{1},t_{2}$. We define the problem of semantic change identification as assigning a value ($y$) to this word,
$f_{\mathcal{C}}(\mathfrak{R}(w,t_{1}),\mathfrak{R}(w,t_{2}))\rightarrow y.$ (9)
Schlechtweg and im Walde, (2020) proposes two notions of semantic change, the binary classification, e.g., $y\in\{0,1\}$, and the graded (continuous) semantic change, e.g., $y\in\mathbf{R}$, where $y$ is a distance between word representations [^6] While in Schlechtweg and im Walde, (2020) they consider the graded change for a particular purpose, i.e., the Jasen-Shannon distance, we generalize $y$ to any distance measure applicable to the representation of words..
The binary classification task is aligned to the classical LSC literature (Blank,, 2012, p. 113) and to the human perception which in many cases are influenced by a sensory threshold (Smith,, 2008, p. 34). This notion of change is interested on the gain/lost of senses in a word. There’s no general threshold criteria for the binary setting, it is not enough to have different representations, but the magnitude of this difference should be aligned to human judgment. For example, in Schlechtweg et al., (2020) they define a criteria for the graph representation based on corpus size. However, there’s no measure on how it correlates with human perception of change and how it compares to other representations.
The graded task is originated from the computational linguistics literature, as in computers a natural way to compare objects is measuring the distance between its representations. This approach for semantic change allows a more in depth analysis of change as the grade of change can be compared across a wide range of contexts (history, cultures, etc.).
@@ 5.6 Semantic Change Characterization
Deriving from the reviewed literature, we define semantic change characterization as a classification problem where we need to assign identified semantic changes to categories, such as $f_{\mathcal{D}},f_{\mathcal{R}},f_{\mathcal{O}}$,
@@
While the first equation only measured some notion of distance between the representations, for characterization we first apply the type function on the representation, to then compare. For example in Buechel et al., (2016) we have a orientation characterization, e.g.,
$f(\mathcal{O}(\mathfrak{R}(w,t_{1})),\mathcal{O}(\mathfrak{R}(w,t_{2})))\rightarrow y,$ (11)
where the $\mathcal{O}$ function is the Valence-Arosal-Domance framework and they measure the distance $f$ through the $\beta$-coefficient of a linear regression.
Again we set two possibilities of change, binary and graded. In this setting the binary form allow us to answer interesting questions in an automatic manner, for example, in which point in time the word ‘head’ started being used in a figurative manner? Another possibility is to track the precise corpus were words pejorated, and then manually understand the socio-linguistic process to this phenomena, saving many hours of historiography work.
@@ 6.1 Dimension
The majority of the studies found during this survey propose methods and tools to characterize how words change in dimension, while the other poles were less studied. Change in dimension seems to be easier to detect than the other categories because it does not rely on subjective interpretation, such as relation or feeling, but mainly on countable quantity (see Section 5). Works along this axis often rely on word sense induction, a research area with well developed tooling and many data sources, such as dictionaries and synsets, used to implement supervised learning.
A major characteristic that differentiate the approaches is the word representation method. We identified four main types of word representation methods, used to study dimension changes in LSC. These methods have different trade-offs in implementation cost, representation power, and interpretability. It is difficult to compare the effectiveness of each method for this state-of-the-art, given the qualitative and exploratory nature of the reported results, the use of different corpora (not always available), and the need for external knowledge to acquire meaning (difficult to reproduce). However, the analysis reveals that word frequency and embeddings-based methods produce better results to identify neologism (new words), while topics and graph-based approaches are more effective at detecting sememes (new senses) over time. A limitation for the approaches implemented based on dictionaries is that the meaning of new words are not always in the dictionaries because there is a temporal delay between the first occurrences of these neologisms and their publication in dictionaries. Another limitation comes with the presence, in dictionaries, of metaphorical senses at the same level as the literal ones, increasing the complexity to make the distinction between ‘dimension’ and ‘relation’. Some authors prefer to group these two poles into one, as in (Tang et al.,, 2013) . The limitation for topics- and graph-based approaches is that the interpretation of the meaning of a topic (or a branch of a graph) is not always clear and can require the intervention of an expert to, for instance, to regroup topics/branches with the same meaning and/or to map these groups to word’s senses.
@@ 6.2 Relation
Outside of this LSC survey, there is a vast body of work proposing methods and tools to predict if a word in a sentence has a metaphorical sense (Heintz et al.,, 2013; Shutova and Sun,, 2013; Hovy et al.,, 2013). These approaches analyze texts, but they do not consider the temporal information, thus they cannot express if a word suffers from metaphorization over time. Some works claim to detect metaphorization of words (e.g., (Tredici et al.,, 2016) ), but they lay on the general hypothesis that changing sense of words implies in a metaphorical change, which still needs to be demonstrated.
For metonymization, we could not find studies that performed metonymy identification in a comparative setting. However, related recent studies investigate the usage of metonymy within different methods to build the material relation identification (Nastase et al.,, 2012; Soares et al.,, 2019; Mathews and Strube,, 2021). We expect that future investigations can extend these approaches to allow detecting metonymization of words.
@@
(Tang et al.,, 2013) mainly rely on dictionaries of heuristics to infer metaphoricity, which served for qualitative analysis, but still require more extensive validation. A single work used ground-truth annotations for metaphors (Schlechtweg et al.,, 2017), but did not reported a final precision of the proposed methodology.
@@ 6.3 Orientation
Although there is a huge body of work proposing solutions for analysing sentiment at sentence level, fewer of them try to infer sentiments at a word-level (Guo et al.,, 2021), and in the word-level, only a small fraction is performed on a comparative manner, as we accessed in this survey.
An important outcome of this survey is that there is an harmonization on how change in feeling are defined and applied. However, we could not observe that there is also an harmonization on how the error analysis for the feeling function is implemented. Some works adopt binary orientation function (positive, negative), while other multidimensional orientations (valence, arousal, dominance) (Buechel et al.,, 2016). There is still further study to analyze which function ($\mathfrak{f}$) is more suitable for characterizing change.
One can observe some converging points from the many approaches analyzed in this survey. This is the case, for instance, for the difficulty to evaluate and compare the obtained results because, for most of the tasks, no gold standard exists. Many papers have raised this problem (Schlechtweg et al.,, 2017), but some initiatives to create reference datasets are under development.